# UniGame: Turning a Unified Multimodal Model Into Its Own Adversary

### v2

@@ 4.2 Consistency Evaluation
We evaluated consistency on two benchmarks: UnifiedBench [yan2025can] and WISE score [niu2025wise]. UnifiedBench is a reconstruction-based benchmark tailored for UMMs, where the “caption–generate–compare” protocol measures how consistently information is preserved when images is converted into text and decoded back into images; the unified score is defined as the similarity between the ground truth and reconstructed images. WISE is a world-knowledge-informed text-to-image benchmark with 1000 knowledge-intensive prompts that are scored by a multimodal judge along consistency with the prompt, realism, and aesthetic quality; its overall Score emphasizes how faithfully the generated image reflects the textual description, making it a natural testbed for text-to-image consistency.
We followed Protocol 1 from [yan2025can] to evaluate the consistency of UniGame. Specifically, we randomly sampled $100$ images from LAION-5B [schuhmann2022laion] as a test set. For each image, the model first generated a caption (understanding), based on which the model then reconstructed the image (generation). Finally, we computed the similarity scores between the original and reconstructed images using four vision-language backbones (CLIP [radford2021learning], SigLIP [zhai2023sigmoid], DINO-v2 [oquab2023dinov2], and DreamSim [fu2023dreamsim]).
@@
TokenFlow-XL [qu2025tokenflow] 14B $77.6$ $43.2$ $76.8$ $87.8$ $71.3$
BAGEL [deng2025emerging] 14B $55.3$ $85.0$
UniWorld-V1 [lin2025uniworld] 12B $58.6$ $83.5$
BLIP3-o [chen2025blip3] 8B $83.1$ $50.6$ $83.5$
Emu3 [wang2024emu3] 8B $75.1$ $31.6$ $58.5$ $85.2$ $62.6$
SEED-X [ge2024seed] 7B $71.2$ $35.6$ $70.1$ $84.1$ $65.2$
Chameleon [team2024chameleon] 7B $66.0$ $22.4$
OminiGen2 [xiao2025omnigen] 7B $53.1$ $79.1$
Liquid [wu2024liquid] 7B $63.5$ $76.8$
Janus-Pro [chen2025janus] 7B $78.2$ $41.0$ $79.2$ $87.4$ $71.4$
Show-o [xie2024show] 1.3B $69.4$ $26.7$ $80.0$
@@
Table 3: Results of text-to-image generation on GenEval [ghosh2023geneval] ($\uparrow$).
@@
SEED-X [ge2024seed] $0.97$ $0.58$ $0.26$ $0.80$ $0.19$ $0.14$ $0.49$
Show-o [xie2024show] $0.95$ $0.52$ $0.49$ $0.82$ $0.11$ $0.28$ $0.53$
D -DiT [li2025dual] $0.97$ $0.80$ $0.54$ $0.76$ $0.32$ $0.50$ $0.65$
TokenFlow-XL [qu2025tokenflow] $0.95$ $0.60$ $0.41$ $0.81$ $0.16$ $0.24$ $0.55$
Chameleon [team2024chameleon] $0.39$
OminiGen2 [xiao2025omnigen] $0.99$ $0.86$ $0.64$ $0.85$ $0.31$ $0.55$ $0.70$
Janus-Pro [chen2025janus] $0.99$ $0.89$ $0.59$ $0.90$ $0.79$ $0.66$ $0.80$
@@ 4.3 Benchmark Results
UniGame improves both understanding and generation. Table 2 and Table 3 show the results on understanding and generation benchmarks, respectively. The results demonstrate significant improvements of UniGame over most competitors in both tasks. Specifically for understanding tasks, UniGame shows an average improvement of $3.1\%$ over SFT, and $3.6\%$ over the baseline model. UniGame also outperforms other larger models such as TokenFlow-XL, Emu3, and BLIP-3o, demonstrating that creating adversarial examples can serve as an effective augmentation approach. As for generation, UniGame outperforms most competitors such as TokenFlow-XL and OminiGen, achieving stronger performance than the base model and SFT. Note that UniGame performs slightly worse than UAE and RecA ($0.82$ vs. $0.86$), which is mainly due to UAE and RecA’s explicit post-training on generation tasks, while ours was primarily trained on understanding tasks.
@@
Table 4: Ablation: UniGame vs embedding perturbation. All methods use the same perturbation budget ($\varepsilon_{\mathrm{max}}{=}0.02$), update schedule, and 16k training steps. We report VQAv2 accuracy (%).
@@
   Full (+ CLIP + Buffer) 83.4
@@ 4.4 Ablation Study
This ablation reveals three key insights: (i) the embedding-level perturbations can only leverage weak adversarial signals ($0.7\%$) because they operate in an abstract space disconnected from visual semantics; (ii) decoder constraints enforce on-manifold perturbations, yielding stronger adversarial training ($2.0\%$); (iii) using CLIP to maintain semantics further amplifies gains by ensuring adversarial samples remain semantically consistent with the query text. Together, these components establish a principled framework for self-adversarial training in UMMs.
@@ 4.6 Extensibility and Efficiency
UniGame remains agnostic to UMM architectures and is computationally efficient compared to other post-training methods. In this section, we evaluate its generality and efficiency using the full set of VQAv2 on 2×H100 (80 GB) with mixed precision. Unless noted otherwise, we use image generation size=$384$ and a global batch size of $8$.
@@
understanding generation consistency
@@
RecA + Ours 36.2 (+0.5) 0.86 () 68.21 (+1.27)
Extensibility. We first implement UniGame as a complement to RecA [xie2025reconstruction]. The results in Table 5 show consistent performance gains. Specifically, UniGame outperforms the original RecA by $0.5$ on MMMU for understanding and $1.27$ on UnifiedBench for consistency, while remaining the same on GenEval. These results indicate that UniGame can serve as a lightweight, plug-and-play post-training module that can be integrated into existing pipelines, requiring only minimal additional computation.
In addition to RecA, we further constructed two architectures: (1) UMM-1 uses a Qwen2.5-VL [yang2025qwen3] backbone with a SigLIP2 [zhai2023sigmoid] understanding encoder and a Stable Diffusion-1.5 [Rombach_2022_CVPR] image branch; (2) UMM-2 keeps the vision/generation stack unchanged and replaces the backbone with GPT-OSS [agarwal2025gpt]. Since GPT-OSS is designed for texts, we inserted a trainable 2-layer MLP that projects vision embeddings into the language space. Table 6 demonstrates that UniGame is agnostic to model architectures and can improve the performance of all three different backbones. Moreover, the gains are achieved with fewer trainable parameters (e.g., $\sim$0.45% for UMM-2 and $\sim$1.43% for UMM-1), indicating its parameter-efficient generalization.
@@
UMM-1 (Qwen2.5-VL) 60.4 66.4 (+6.0%) $\sim$1.43% (100.3M / 7B)
UMM-2 (GPT-OSS) 28.9 53.2 (+24.3%) $\sim$0.45% (133.9M / 30B)
@@
ReCA 34.7 35.7 (+1.0%) $\sim$91% (1.4B / 1.5B)
UAE $\sim$1% (0.1B / 11B)
UniGame 41.0 43.8 (+2.8%) $\sim$1% (100.3M / 7B)
Efficiency. We further evaluate the efficiency of UniGame in comparison with RecA [xie2025reconstruction] and UAE [yan2025can] on MMMU. Table 7 shows that while achieving stronger performance, UniGame uses fewer trainable parameters, indicating its efficiency over existing post-training approaches.
@@ 4.7 Convergence and Hyperparameter Analysis
Finally, we present a large-scale analysis on the convergence and hyperparameters (e.g., the hard buffer threshold $\tau$, trade-off $\beta$, and perturbation budget $\delta$; see Figure 11). To study the training dynamic, we also conduct extensive ablation study on the learning rates of two major minmax opponents in Figure 6(b). Further, we systematically study the minimax dynamics by visualizing the optimization trajectory of each run. The best configuration yields a well-behaved minimax trajectory, where the two players alternate smoothly without divergence (see Appendix 9). We also probe the self-play dynamics between the two opponents and clearly observe the interaction: the two branches alternately dominate the training objective, exhibiting a stable tug-of-war behavior, and change of dominance, see Figure 10 and 12. More details in Appendix F demonstrate that UniGame offers a steady training process and stays relatively robust to different hyperparameter choices. Appendix G further presents some theoretical insights.
@@ 5 Conclusion and Limitation
Appendix
A Algorithm Details ........................................................................................................................................................................A
B Training Details ........................................................................................................................................................................B
C Detailed Experimental Results ........................................................................................................................................................................C
D Robustness Results ........................................................................................................................................................................D
E Details on Case Study ........................................................................................................................................................................E
F Convergence and Hyperparameter Analysis ........................................................................................................................................................................F
G Theoretical Insights ........................................................................................................................................................................G
@@ B.1 Training and Testing Data
Data volumes. Unless otherwise noted, we follow the official training/evaluation splits and report results on the standard benchmarks. Training uses VQAV2 train-split [balanced_vqa_v2] is a large-scale visual question answering benchmark (hundreds of thousands of image–question pairs) collected from MS-COCO images with crowd-sourced free-form answers; it emphasizes grounded visual reasoning under natural images. CC3M [sharma2018conceptual] (training only) is a large web-scale image–caption corpus ($\sim$3M pairs in the full set); we use a filtered subset of 100k as text–image supervision for the generative branch.
@@
  • VQAv2 test-dev [balanced_vqa_v2]: the official VQAv2 test-dev split contains 104 000104\,000 questions; evaluation is via the online server. [^7] Counts from the official VQA site; see also recent reports confirming 104K for test-dev.
  • MMMU [yue2024mmmu]: a college-level, multi-discipline benchmark with 11 50011\,500 questions in total (we report on the official test set).
  • POPE [li2023evaluating]: object-hallucination evaluation with a balanced, image-grounded design; the test split has $9000$ QA pairs.
  • MMBench [liu2024mmbench]: curated multiple-choice suite; dev $1164$ and test $1784$ questions (4:6 split of $\sim$3K).
  • GenEval [ghosh2023geneval]: object/layout/attribute–focused T2I evaluation with $553$ prompts (reference-free automatic checks).
@@
  • WISE [niu2025wise]: knowledge-informed T2I evaluation with $1000$ structured prompts across 25 subdomains.
@@
  • AdVQA [li2021adversarial]: human-in-the-loop adversarial VQA; total size reported as $\sim$46 80746\,807 examples (commonly used splits include $\sim$$5123$ val / $\sim$23 39923\,399 test).
@@ B.4 Hard Samples
UniGame added a hard sampler buffer to select only the challenging adversarial samples for training. Figure 6(a) shows some challenging examples in our experiments.
@@ Appendix C Detailed Experimental Results
C.1 Learning Rate Ratio Ablation
@@
C.2 Motivation Experiments
@@
(a) Cases drawn from the hard-sample buffer that successfully challenged the model.
(b) VQA accuracy across different adversarial ratios, with the best reaching 83.4%.
Figure 6: Hard examples and learning-rate ratio ablation. (a) Representative hard cases mined into the replay buffer. (b) Training dynamics under different adversarial update ratios.
Figure 7: Perturbation Sweetspot
@@ G.1 Convergence of the Minimax Self-Play Dynamics
$\nabla_{\theta_{U}}\mathcal{L}(\theta_{U}^{\ast},\theta_{C}^{\ast})=0,\qquad\nabla_{\theta_{C}}\mathcal{L}(\theta_{U}^{\ast},\theta_{C}^{\ast})=0.$
@@ G.3 Manifold-Expanding Effect of Decoder-Constrained Perturbations
$\tilde{x}=G(z+\delta),\quad\tilde{x}\in\mathcal{M}$, (15)
@@ G.4 Summary of Theoretical Insights
WARNING: do not forget to delete the supplementary pages from your submission

### v3

@@ 4.2 Consistency Evaluation
We evaluated consistency on two benchmarks: UnifiedBench [43] and WISE score [23]. UnifiedBench is a reconstruction-based benchmark tailored for UMMs, where the “caption–generate–compare” protocol measures how consistently information is preserved when images is converted into text and decoded back into images; the unified score is defined as the similarity between the ground truth and reconstructed images. WISE is a world-knowledge-informed text-to-image benchmark with 1000 knowledge-intensive prompts that are scored by a multimodal judge along consistency with the prompt, realism, and aesthetic quality; its overall Score emphasizes how faithfully the generated image reflects the textual description, making it a natural testbed for text-to-image consistency.
We followed Protocol 1 from [43] to evaluate the consistency of UniGame. Specifically, we randomly sampled $100$ images from LAION-5B [34] as a test set. For each image, the model first generated a caption (understanding), based on which the model then reconstructed the image (generation). Finally, we computed the similarity scores between the original and reconstructed images using four vision-language backbones (CLIP [29], SigLIP [46], DINO-v2 [27], and DreamSim [5]).
@@
TokenFlow-XL [28] 14B $77.6$ $43.2$ $76.8$ $87.8$ $71.3$
BAGEL [4] 14B $55.3$ $85.0$
UniWorld-V1 [19] 12B $58.6$ $83.5$
BLIP3-o [1] 8B $83.1$ $50.6$ $83.5$
Emu3 [37] 8B $75.1$ $31.6$ $58.5$ $85.2$ $62.6$
SEED-X [7] 7B $71.2$ $35.6$ $70.1$ $84.1$ $65.2$
Chameleon [36] 7B $66.0$ $22.4$
OminiGen2 [41] 7B $53.1$ $79.1$
Liquid [39] 7B $63.5$ $76.8$
Janus-Pro [2] 7B $78.2$ $41.0$ $79.2$ $87.4$ $71.4$
Show-o [42] 1.3B $69.4$ $26.7$ $80.0$
@@
Table 3: Results of text-to-image generation on GenEval [8] ($\uparrow$).
@@
SEED-X [7] $0.97$ $0.58$ $0.26$ $0.80$ $0.19$ $0.14$ $0.49$
Show-o [42] $0.95$ $0.52$ $0.49$ $0.82$ $0.11$ $0.28$ $0.53$
D -DiT [18] $0.97$ $0.80$ $0.54$ $0.76$ $0.32$ $0.50$ $0.65$
TokenFlow-XL [28] $0.95$ $0.60$ $0.41$ $0.81$ $0.16$ $0.24$ $0.55$
Chameleon [36] $0.39$
OminiGen2 [41] $0.99$ $0.86$ $0.64$ $0.85$ $0.31$ $0.55$ $0.70$
Janus-Pro [2] $0.99$ $0.89$ $0.59$ $0.90$ $0.79$ $0.66$ $0.80$
@@ 4.3 Benchmark Results
UniGame improves both understanding and generation. Table 2 and 3 show the results on understanding and generation benchmarks, respectively. The results demonstrate significant improvements of UniGame over most competitors in both tasks. Specifically for understanding tasks, UniGame shows an average improvement of $3.1\%$ over SFT, and $3.6\%$ over the baseline model. UniGame also outperforms other larger models such as TokenFlow-XL, Emu3, and BLIP-3o, demonstrating that creating adversarial examples can serve as an effective augmentation approach. As for generation, UniGame outperforms most competitors such as TokenFlow-XL and OminiGen, achieving stronger performance than the base model and SFT. Note that UniGame performs slightly worse than UAE and RecA ($0.82$ vs. $0.86$), which is mainly due to UAE and RecA’s explicit post-training on generation tasks, while ours was primarily trained on understanding tasks.
@@
Table 4: Ablation: UniGame vs embedding perturbation. All methods use the same perturbation budget ($\varepsilon_{\mathrm{max}}{=}0.02$), update schedule, and training steps (16k). For embedding-only baselines, we apply perturbations directly in the visual token space without decoding. We report VQAv2 accuracy (%).
@@
Full (+ CLIP + Buffer) $\mathbf{83.4}$
@@ 4.4 Ablation Study
We further ablate the perturber and hard-sample buffer capacity. For the perturber, a 3-layer MLP (83.4%) outperforms both the 2-layer variant (82.8%) and the deeper 4-layer variant (81.2%), indicating that moderate capacity best balances expressiveness. For the buffer, a size of 50 yields the best accuracy (83.4%), while smaller sizes of 30 (83.1%) and 10 (82.5%) provide insufficient diversity. Key insights: (i) the embedding-level perturbations can only leverage weak adversarial signals ($0.7\%$) because they operate in an abstract space disconnected from visual semantics; (ii) decoder constraints enforce on-manifold perturbations, yielding stronger adversarial training ($2.0\%$); (iii) using CLIP to maintain semantics further amplifies gains by ensuring adversarial samples remain semantically consistent with the query text. Together, these components establish a principled framework for self-play training in UMMs.
@@ 4.6 Extensibility and Efficiency
UniGame remains agnostic to UMM architectures and is computationally efficient compared to other post-training methods. In this section, we evaluate its generality and efficiency using the full set of VQAv2 on 2×H100 (80 GB) with mixed precision. Unless noted otherwise, we use image generation size $384$ and a global batch size of $8$.
@@
understanding generation ocnsistency
@@
RecA + Ours $36.2$ (+0.5+0.5) $0.86$ (--) $68.21$ (+1.27+1.27)
Extensibility. We implement UniGame as a complement to RecA [38]. Table 5 shows consistent performance gains. method outperforms the RecA by $0.5$ on MMMU for understanding and $1.27$ on UnifiedBench for consistency, while remaining the same on GenEval. These results indicate that UniGame can serve as a lightweight, plug-and-play post-training module that can be integrated into existing pipelines, requiring only minimal additional computation.
In addition to RecA, we further constructed two architectures: (1) UMM-1 uses a Qwen2.5-VL [44] backbone with a SigLIP2 [46] understanding encoder and a Stable Diffusion-1.5 [33] image branch; (2) UMM-2 keeps the vision/generation stack unchanged and replaces the backbone with GPT-OSS [26]. Since GPT-OSS is designed for texts, we inserted a trainable 2-layer MLP that projects vision embeddings into the language space. We further applied UniGame to two backbones: BLIP-3o [1] and Chameleon [36]. All show consistent gains (Appendix C.1). Table 6 shows that UniGame is agnostic to model architectures and can improve the performance of different backbones. Moreover, the gains are achieved with fewer trainable parameters (e.g., $\sim$0.45% for UMM-2 and $\sim$1.43% for UMM-1), indicating its parameter-efficient generalization.
@@
[…]
Efficiency. We further evaluate the efficiency of UniGame in comparison with RecA [38] and UAE [43] on MMMU. Table 7 shows that while achieving stronger performance, UniGame uses fewer trainable parameters, indicating its efficiency over existing post-training approaches.
@@
[…]
UAE  $\sim\,1\%~(0.1\mathrm{B}/11\mathrm{B})$
[…]
@@ 4.7 Convergence and Hyperparameter Analysis
Finally, we present a large-scale analysis on the convergence and hyperparameters (e.g., the hard buffer threshold $\tau$, trade-off $\beta$, and perturbation budget $\delta$; see Figure 12). To study the training dynamic, we also conduct extensive ablation study on the learning rates of two major minmax opponents in Figure 7. Further, we systematically study the minimax dynamics by visualizing the optimization trajectory of each run. The best configuration yields a well-behaved minimax trajectory, where the two players alternate smoothly without divergence, see Appendix 10. We also probe the self-play dynamics between the two opponents and clearly observe the interaction: the two branches alternately dominate the training objective, exhibiting a stable tug-of-war behavior, and change of dominance, see Figure 11 and 13. More details in Appendix F demonstrate that UniGame offers a steady training process and stays relatively robust to different hyperparameter choices. Appendix G further presents some theoretical insights.
@@ 5 Conclusion and Limitation
Acknowledgments
This paper is partially supported by unrestricted gift from Google, William & Mary Faculty Research Award, and Modal Academic Compute Award. The authors acknowledge William & Mary Research Computing for providing computational resources and/or technical support that have contributed to the results reported within this paper. URL: https://www.wm.edu/it/rc.
Supplementary Material
Appendix Contents
A   Algorithm Details ....................................................................................................................A
B   Training Details ....................................................................................................................B
C   Detailed Experimental Results ....................................................................................................................C
D   Robustness Results ....................................................................................................................D
E   Details on Case Study ....................................................................................................................E
F   Convergence and Hyperparameter Analysis ....................................................................................................................F
G   Theoretical Insights ....................................................................................................................G
@@ B.1 Training and Testing Data
Data volumes. Unless otherwise noted, we follow the official training/evaluation splits and report results on the standard benchmarks. Training uses VQAV2 train-split [11] is a large-scale visual question answering benchmark (hundreds of thousands of image–question pairs) collected from MS-COCO images with crowd-sourced free-form answers; it emphasizes grounded visual reasoning under natural images. CC3M [35] (training only) is a large web-scale image–caption corpus ($\sim$3M pairs in the full set); we use a filtered subset of 100k as text–image supervision for the generative branch.
@@
  • VQAv2 test-dev [11]: the official VQAv2 test-dev split contains 104 000104\,000 questions; evaluation is via the online server. [^7] Counts from the official VQA site; see also recent reports confirming 104K for test-dev.
  • MMMU [45]: a college-level, multi-discipline benchmark with 11 50011\,500 questions in total (we report on the official test set).
  • POPE [17]: object-hallucination evaluation with a balanced, image-grounded design; the test split has $9000$ QA pairs.
  • MMBench [20]: curated multiple-choice suite; dev $1164$ and test $1784$ questions (4:6 split of $\sim$3K).
  • GenEval [8]: object/layout/attribute–focused T2I evaluation with $553$ prompts (reference-free automatic checks).
@@
  • WISE [23]: knowledge-informed T2I evaluation with $1000$ structured prompts across 25 subdomains.
@@
  • AdVQA [16]: human-in-the-loop adversarial VQA; total size reported as $\sim$46 80746\,807 examples (commonly used splits include $\sim$$5123$ val / $\sim$23 39923\,399 test).
@@ B.4 Hard Samples
UniGame added a hard sampler buffer to select only the challenging adversarial samples for training. Figure 6 shows some challenging examples in our experiments.
Figure 6: Cases are drawn from the hard-sample buffer and represent failure cases that successfully challenged the model.
@@ Appendix C Detailed Experimental Results
C.1 Additional Backbone Validation
To validate the generality of UniGame beyond Janus-Pro-7B, we applied it to two additional UMM backbones: BLIP-3o [1] (diffusion-based) and Chameleon [36] (auto-regressive), both trained on a 0.4 split of the VQAv2 training set under-matched settings. All backbones show positive consistency gains: BLIP-3o (+2.5, MMMU 51.2, GenEval 0.54), and Chameleon (+2.2, VQAv2 40.2 with +1.7 gain, MMMU 24.0, GenEval 0.40). These results confirm that UniGame generalizes across different UMM architectures, including both auto-regressive and diffusion-based designs.
C.2 Learning Rate Ratio Ablation
Figure 7: Training dynamics analysis. VQA accuracy evolution across different adversarial ratios, with best achieving optimal performance at 83.4%.
@@ C.2 Learning Rate Ratio Ablation
C.3 Motivation Experiments
Figure 9: Perturbation Sweetspot
@@ G.1 Convergence of the Minimax Self-Play Dynamics
$\nabla_{\theta_{U}}\mathcal{L}(\theta_{U}^{\ast},\theta_{C}^{\ast})=0,\hskip 28.80008pt\nabla_{\theta_{C}}\mathcal{L}(\theta_{U}^{\ast},\theta_{C}^{\ast})=0.$
@@ G.3 Manifold-Expanding Effect of Decoder-Constrained Perturbations
$\tilde{x}=G(z+\delta),\qquad\tilde{x}\in\mathcal{M}$, (15)